# Momentum dependent quantum Ruelle-Pollicott resonances in translationally invariant many-body systems

### v2

@@ II.1 Truncated propagator of translated local operators
which is from the sector with $k=\pi$. Also, by combining different $k$ sectors one can as well treat non-extensive operators, for instance, a strictly local ones.
@@ II.2 Some properties of $M(k)$
Let us mention at this point that the properties of matrix $M(k)$ are rather interesting and warrant future studies. As already noted in Ref. Prosen when studying the $k=0$ case $M$ has a self-similar structure (which is also reflected in a self-similar structure of eigenvectors). Here we just list few numerical observations that will help us understand behavior of RP resonances and the decay of correlations.
We start with singular values of $M(k)$ that turn out to be independent of $k$ as well as of $h_{\rm x}$ and $h_{\rm z}$. They depend only on $\tau$. In fact, there are only 3 different singular values with simple analytical form,
@@
Independence of singular values on the parameters of single qubit gates is easy to see. Namely, the unitary propagator (I.3) is a product of single qubit gates and two qubit gates. Because the truncation we use is a truncation in the support of operators, and single qubit gates do not change the support, the truncated propagator can be written as a product of two matrices,
@@
where $M_{\rm zz}$ is non-unitary and corresponds to transformation by all two qubit gates $U_{\rm zz}={\rm e}^{{\rm i}\tau\sum_{j}\sigma^{\rm z}_{j}\sigma^{\rm z}_{j=1}}$, while $O$ is orthogonal matrix and describes the action of the remaining single qubit gates. Eq.(16) immediately means that $M$ and $M_{\rm zz}$ have the same singular values (though different eigenvalues!). Therefore, the singular values of the full $M$ depend only on $\tau$. Their exact functional form should presumably follow from simple form of $M_{\rm zz}$.
Moving to eigenvalues, as we shall see later (Fig. 10), the bulk of the spectrum of eigenvalues of $M(k)$ will consist of an annular cloud of eigenvalues. Such an annular region of eigenvalues is for instance obtained foot3 in a singular value decomposition motivated random matrix ensemble of form $U\,s\,V$, where $U$ and $V$ are Haar random unitary while $s$ is a fixed real diagonal matrix (of e.g. singular values) zee01 ; ring . Nice single-ring theorem also says zee01 that one can express the inner radius $r_{\rm in}$ and the outer one $r_{\rm out}$ of the annulus in terms of averages over singular values $s_{j}$,
@@
where $r_{\rm out}$ is given by Eq.(18). The bound in particular limits the smallest possible leading RP regardless of parameters to be $|\lambda_{1}|>{\rm min}(r_{\rm out})=\sqrt{5/12}\approx 0.64$ (see also brown curve in Fig. 2). In the cases we tested the bound (19) is always satisfied foot2 .
From singular values (14) we see that $\tau=\pi/4$ is a special point (as well as any odd multiple of $\pi/4$). This is the point where the interaction gate of the KI becomes a signed permutation matrix (Appendix B), and the truncated propagator becomes singular – there are only $5\cdot 4^{r-2}$ nonzero singular values that are all $1$ with the rest being zero. This is also the number of nonzero eigenvalues of $M(k)$ at $\tau=\pi/4$. If one in addition sets $h_{\rm x}=1$, i.e., a one parameter family of KI models with $\tau=\pi/4$ (or an odd multiple of $\pi/4$) and $h_{\rm x}=1$ one has an even more special situation with the unitary propagator $U$ being dual unitary DU . For the dual unitary case we note that while the number of nonzero singular values stays the same $5\cdot 4^{r-2}$, the number of nonzero eigenvalues of $M(k)$ further decreases to just $2^{r-1}$. They are all independent of $k$ (in-line with a spatial delta function dependence of correlations which are for dual unitary circuits nonzero only on a light-cone), though they do depend on $h_{\rm z}$. If one also sets $h_{\rm z}=1$ the KI model circuit is dual unitary as well as Clifford, and all eigenvalues of $M(k)$ become 00. These interesting properties suggest further study of the case $\tau=\pi/4$.
@@ II.3 Kicked Ising resonances
Let us now numerically calculate foot1 the leading eigenvalues of the matrix $M(k)$ (10) for the KI model (I.3). We pick parameters similar to those used in Ref. Mori , namely $h_{\rm x}=0.9$ and $h_{\rm z}=0.8$, and two values of $\tau=0.65$ and $\tau=0.75$. In Fig. 1 we show the $k$-dependence of the eigenvalues $\lambda_{1}$ with the largest modulus.
Figure 1: Dependence of the leading RP eigenvalue on $k$, all for $h_{x}=0.9$, $h_{z}=0.8$. Full squares show the values in the limit $r\to\infty$ for $k=0$ (upper square for $0^{+}$, lower for $0^{-}$) and $k=\pi$. Lighter colored curves (cyan for $\tau=0.65$ and orange for $\tau=0.75$) show results for $r=9$ indicating how fast $|\lambda_{1}|$ converges with $r$. Small points along the curves indicate values of $k$ where $\lambda_{1}$ is real (for $\tau=0.65$ this are most of $k$, for $\tau=0.75$ only very small ones before the first kink).
We can see several interesting features. First, the largest eigenvalue comes from the $k=0^{+}$ sector. The $k$ dependence is expectedly nontrivial, with several kinks visible at values where one has a collision of eigenvalues.
Even more interesting and informative is the dependence of the leading resonance on $\tau$. In Fig. 2 we focus on three different momentum sectors, $k=0^{\pm}$ and $k=\pi$, in which the leading RP resonance is expected to give the decay rate of the autocorrelation function of e.g. magnetization $Z$ ($k=0^{+}$), current $J$ ($k=0^{-}$), and staggered magnetization $sZ$ ($k=\pi$). There are characteristic dips around $\tau\approx 0.8$ and $\tau\approx 2.5$, where one therefore expects the strongest chaoticity, i.e., the fastest decay of correlations. Overall, we again see that in $k=0^{+}$ one has the largest leading RP resonance. Looking only at the $k=0^{+}$ we see (as was noted already in Ref. Prosen ; Prosen07 ) that there are large section of $\tau$ values where $|\lambda_{1}|$ appears to be essentially $1$, which would suggest that correlations do not decay, i.e., as if one would have an almost “integrable” system (rigorous integrability is of course highly unlikely). Note that we do not have in mind the interval $1.5\lesssim\tau\lesssim 2.0$ where one indeed has three quite close integrable points at $\tau=\pi/2$ (integrability type (iv) from introduction), and $\tau=\pi/(2h_{\rm x})$ and $\tau=\pi/(2h_{\rm z})$ (type (iii) integrability), but regions outside of this interval.
@@
However, upon inspecting other sectors we see that things are different. The values of $|\lambda_{1}|$ there are smaller than $1$, suggesting decay of correlations. We shall explore in more details all those situations for the three indicated selected values of $\tau$ in the following sections. For now let us just say that they correspond to different dynamical regimes, e.g., chaotic with exponential mixing vs. just mixing, and that having access to those other momentum sectors is instrumental in determining the correct dynamical behavior.
Finally, let us comment on the convergence of $\lambda_{1}$ with $r$. We can see in Fig. 2 that the speed of convergence varies very much with $\tau$ as well as with $k$. For instance, in the $k=0^{+}$ sector $|\lambda_{1}|$ has converged already at $r=10$ for almost all shown values of $\tau$ (red for $r=10$ and orange curve for $r=8$ are indistinguishable except in a very narrow region around the minimum at $\tau\approx 0.8$). For $k=0^{-}$ convergence is good for $\tau>0.65$, but much worse for smaller $\tau$. Situation is least converged for $k=\pi$ (note also that we show only $r=8$), however there is again strong asymmetry; the eigenvalue is converged for larger $\tau$, while not at smaller $\tau$. Slow convergence at very small $\tau$ can likely be traced back to the fact that all eigenvalues are very close to $1$, e.g., both inner and outer radius (18) of a ring that contains most of eigenvalues converge to $1$ as $\tau\to 0$. For $\tau$ that is an even multiple of $\pi/4$ (the dips in Fig. 2) the inner radius $r_{\rm in}$ instead becomes 00, however the number of nonzero eigenvalues of $M(k)$ is smaller and the matrix is strongly non-normal, likely contributing to slow convergence with $r$.
The rich dependence on three relevant parameters calls for a detailed investigation of the KI phase diagram. In the present work we will focus only on three representative cases, showing that there are different dynamical regimes in a quantum many body system. In particular, as we shall see, the system is not always chaotic (see also previous results in Ref. Prosen ). That is, even in the thermodynamic limit and for finite breaking of integrability one needs not have chaos. This is very interesting result, akin to what would be in classical system a consequence of the celebrated KAM theorem ott , despite having no quantum analog of the KAM theorem.
@@ III Quantum chaotic regime
Considering that all $|\lambda_{1}(k)|<1$ we can say that we are in a quantum chaotic regime. Therefore one expects that all correlation functions should asymptotically decay to zero exponentially as $\sim|\lambda_{1}|^{t}$, and, considering that the eigenvalue depends on $k$, different operators will decay with a different rate. We show in Fig. 4 that this is indeed the case. We compare numerically calculated $C(t)$ and the predicted asymptotic decay based on the leading RP resonance, always finding agreement (Fig. 3).
@@
decay the slowest. On the other hand, the “current” $J$ (12) decays faster because it is from the odd $k=0^{-}$ sector that has smaller $|\lambda_{1}|$. Still faster is the decay of the staggered magnetization $sZ$ (13), which is from $k=\pi$ ($sZ$ decays in an oscillatory manner therefore we in this case always plot $|C(t)|$). By superimposing operators with different $k$ we can study operators with any spatial dependence. The extreme case is a strictly local operator, for instance polarization $\sigma^{\rm x}_{L/2}$ at the middle size, which is a uniform superposition of all $k$ modes. Because the $k=0^{+}$ contribution will decay the slowest we expect that such local autocorrelation function will asymptotically also decay as $|\lambda_{1}(k=0^{+})|^{t}$. The correlation function, defined in this case without an $1/L$ prefactor,
@@ III.1 Slow convergence with system size
However, looking at the magnetization $Z$ (20) in Fig. 7(a) an exponential decay is seen with a rate that is not given by $|\lambda_{1}(0^{+})|$. This is rather puzzling considering that the RP resonance $\lambda_{1}(0^{+})$ converges rather quickly (Fig. 5), and that this seeming incorrect decay persists over almost 3 decades all the way till finite-size fluctuations become noticeable at very small values of $C(t)$. To understand how can that be we have looked at the evolution of $Z$ by the truncated propagator $M(0^{+})$ whose largest eigenvalue is clearly $\lambda_{1}\approx 0.825$ (and is different that the observed decay that goes as $\sim 0.76^{t}$). Namely, within the truncation to $r$-site operators the corresponding correlation function can be simply written as
@@
where the initial vector $\mathbf{y}$ is one corresponding to the operator $a\in{\cal S}_{r}$, in our case to $a=\sigma^{\rm z}$. In a basis where $a=\sigma^{\rm z}$ is chosen as one basis vector one therefore needs only the appropriate diagonal matrix element of the truncated matrix $M^{t}$.
@@
In Fig. 7(b) we can see that the above truncation approximation to $C(t)$ (Eq.24, curves with small symbols in Fig. 7(b)) does asymptotically decay as $\lambda_{1}^{t}$, but the asymptotic decay start rather late, say around $t>40$ when $C(t)$ is already quite small (because the corresponding expansion coefficient is small). We can also see that, while the $r=6$ approximation does not yet capture the true decay for $t<30$, higher truncations like $r=10$ or $r=12$ do. We can conclude that in order to see the true asymptotic decay $|\lambda_{1}|^{t}$ one needs to look at small values $C(t)<10^{-6}$, and therefore in order to even get a glimpse of it in a single random state evolution one would need large systems with $L>40$.
Such decay where one initially has faster relaxation $\sim 0.76^{t}$ which then eventually does goes into asymptotic $\sim 0.82^{t}$ given by the largest eigenvalue is reminiscent of a recently discovered two-step relaxation in random circuits PRX in which the initial decay is determined by the pseudospectrum rather than the spectrum U4 . It can happen when dealing with non-normal matrices, e.g., non-Hermitian, that are badly conditioned. In such cases the spectrum is a rather singular object that is very sensitive to perturbations while the pseudospectrum trefethen is robust. It seems though that the effect here is different in one crucial aspect: the window in which one has the “incorrect” decay $\sim 0.76^{t}$ is here of fixed width whereas in the mentioned random circuits it diverges with system size. This divergence comes because the condition number of those matrices in random circuits diverges with system size, allowing for increasing window size as one increases the system size. To get an idea on the condition number of our $M(k=0^{+})$ we have looked at its singular values. Namely, for a nonsingular matrix $M$ its condition number is simply $\kappa=s_{\rm max}(M)/s_{\rm min}(M)$, where $s_{\rm max}$ and $s_{\rm min}$ are maximal and minimal singular values. Using exact expressions for singular values (14) we see the the maximal one is $1$, while the minimal one is $\cos^{2}{(2\tau)}$, both independently of $r$ or any other parameters. The matrix $M$ becomes singular only at special $\tau=\pi/4\approx 0.78$. Because our $\tau=0.75$ is very close to that singularity we have a rather large $\kappa\approx 200$, which though is finite and in particular does not increase with $r$. Therefore, there is no true two-step relaxation; the initial decay is just a transient that happens to hold for a rather large range of $C(t)$. Nevertheless, slow convergence of the autocorrelation function towards its asymptotic decay can be traced back to strong non-normality of $M(k)$. An interesting open question is, whether perhaps at $\tau=\pi/4$ one could have some effects of strong non-normality due to some singular values becoming zero.

### v3

@@ II.1 Truncated propagator of translated local operators
which is from the sector with $k=\pi$. Also, by combining different $k$ sectors (Fourier transformation) one can as well treat non-extensive operators, for instance, strictly local ones.
@@ II.2 Some properties of $M(k)$
Let us mention at this point that the properties of matrix $M(k)$ are rather interesting and warrant future studies. As already noted in Ref. Prosen , when studying the $k=0$ case, $M$ has a self-similar structure (which is also reflected in a self-similar structure of eigenvectors). Here we just list few numerical observations that will help us understand the behavior of RP resonances and the decay of correlations.
We start with singular values of $M(k)$ that turn out to be independent of $k$ as well as of $h_{\rm x}$ and $h_{\rm z}$. They depend only on $\tau$. In fact, there are only 3 different singular values having a simple analytical form,
@@
Independence of singular values on the parameters of single qubit gates is easy to see. Namely, the unitary propagator (I.3) is a product of single qubit gates and two qubit gates. Because the truncation we use is a truncation in support of operators, and single qubit gates do not change the support, the truncated propagator can be written as a product of two matrices,
@@
where $M_{\rm zz}$ is non-unitary and corresponds to transformation by all two qubit gates $U_{\rm zz}={\rm e}^{{\rm i}\tau\sum_{j}\sigma^{\rm z}_{j}\sigma^{\rm z}_{j=1}}$, while $O$ is an orthogonal matrix and describes the action of the remaining single qubit gates. Eq.(16) immediately means that $M$ and $M_{\rm zz}$ have the same singular values (though different eigenvalues!). Therefore, the singular values of the full $M$ depend only on $\tau$. Their exact functional form should presumably follow from the simple form of $M_{\rm zz}$.
Moving to eigenvalues, as we shall see later (Fig. 10), the bulk of the spectrum of eigenvalues of $M(k)$ will consist of an annular cloud of eigenvalues. Such an annular region of eigenvalues is for instance obtained foot3 in a singular value decomposition motivated random matrix ensemble of form $U\,s\,V$, where $U$ and $V$ are Haar random unitary while $s$ is a fixed real diagonal matrix (of e.g. singular values) zee01 ; ring . A nice single-ring theorem also says zee01 that one can express the inner radius $r_{\rm in}$ and the outer one $r_{\rm out}$ of the annulus in terms of averages over singular values $s_{j}$,
@@
where $r_{\rm out}$ is given by Eq.(18). The bound in particular limits the smallest possible leading RP resonance, regardless of parameters, to be $|\lambda_{1}|>{\rm min}(r_{\rm out})=\sqrt{5/12}\approx 0.64$ (see also brown curve in Fig. 2). In the cases we tested the bound (19) is always satisfied foot2 .
From singular values (14) we see that $\tau=\pi/4$ is a special point (as well as any odd multiple of $\pi/4$). This is the point where the interaction gate of the KI becomes a signed permutation matrix (Appendix B), and the truncated propagator becomes singular – there are only $5\cdot 4^{r-2}$ nonzero singular values that are all $1$ with the rest being zero. This is also the number of nonzero eigenvalues of $M(k)$ at $\tau=\pi/4$. If one in addition sets $h_{\rm x}=1$, i.e., a one parameter family of KI models with $\tau=\pi/4$ (or an odd multiple of $\pi/4$) and $h_{\rm x}=1$, one has an even more special situation with $U$ being dual unitary DU . For the dual unitary case we note that while the number of nonzero singular values stays the same $5\cdot 4^{r-2}$, the number of nonzero eigenvalues of $M(k)$ further decreases to just $2^{r-1}$. They are all independent of $k$ (in-line with a spatial delta function dependence of correlations which are for dual unitary circuits nonzero only on a light-cone), though they do depend on $h_{\rm z}$. If one also sets $h_{\rm z}=1$ the KI model circuit is dual unitary as well as Clifford, and all eigenvalues of $M(k)$ become 00. These interesting properties suggest further study of the case $\tau=\pi/4$.
Figure 1: Dependence of the leading RP eigenvalue on $k$, all for $h_{x}=0.9$, $h_{z}=0.8$. Full squares show the values in the limit $r\to\infty$ for $k=0$ (upper square for $0^{+}$, lower for $0^{-}$) and $k=\pi$. Lighter colored curves (cyan for $\tau=0.65$ and orange for $\tau=0.75$) show results for $r=9$ indicating how fast $|\lambda_{1}|$ converges with $r$. Small points along the curves indicate values of $k$ where $\lambda_{1}$ is real (for $\tau=0.65$ these are most of $k$, for $\tau=0.75$ only very small ones before the first kink).
Finally, possible eigenvalues $\lambda=1$ of $M(k)$ at a finite truncation $r$ would indicate an exact conserved local operator supported on at most $r$ sites. Because our model is in general chaotic without any conservation laws there are no such eigenvalues except at mentioned special integrable points. Note that the translation operator $T$ can not be written as an exponential of a local generator that would be called a (quasi)momentum operator. While $T$ can be written as a product of nearest-neighbor swap gates $T_{j,j+1}=\exp{[-{\rm i}\frac{\pi}{4}(\sigma^{\rm x}_{j}\sigma^{\rm x}_{j+1}+\sigma^{\rm y}_{j}\sigma^{\rm y}_{j+1}+\sigma^{\rm z}_{j}\sigma^{\rm z}_{j+1}-\mathbbm{1})]}$, trying to write it as $T={\rm e}^{{\rm i}\,\hat{p}}$, the generator $\hat{p}$ is not local. Therefore, we don’t have any eigenvalue $1$ (or close to $1$) of $M(k=0)$ for any finite $r$ that would correspond to the (quasi)momentum $\hat{p}$. In other words, for any finite $r$ the (quasi)momentum operator $\hat{p}$ is not (even close to) conserved. This makes lattice models with our truncation a bit different than continuous systems in an appropriate basis where presumably $\hat{p}\sim-{\rm i}\partial_{x}$ would be included.
@@ II.3 Kicked Ising resonances
Let us now numerically calculate foot1 the leading eigenvalues of the matrix $M(k)$ (10) for the KI model (I.3). We pick parameters similar to those used in Ref. Mori , namely $h_{\rm x}=0.9$ and $h_{\rm z}=0.8$, and two values of $\tau=0.65$ and $\tau=0.75$. In Fig. 1 we show the $k$-dependence of the eigenvalue $\lambda_{1}$ with the largest modulus. We can see several interesting features. First, the largest eigenvalue comes from the $k=0^{+}$ sector. The $k$ dependence is expectedly nontrivial, with several kinks visible at values where one has a collision of eigenvalues.
@@
Even more interesting and informative is the dependence of the leading resonance on $\tau$. In Fig. 2 we focus on three different momentum sectors, $k=0^{\pm}$ and $k=\pi$, in which the leading RP resonance is expected to give the decay rate of the autocorrelation function of e.g. magnetization $Z$ ($k=0^{+}$), current $J$ ($k=0^{-}$), and staggered magnetization $sZ$ ($k=\pi$). There are characteristic dips around $\tau\approx 0.8$ and $\tau\approx 2.5$, where one therefore expects the strongest chaoticity, i.e., the fastest decay of correlations. Overall, we again see that in $k=0^{+}$ one has the largest leading RP resonance. Looking only at the $k=0^{+}$ we see (as was noted already in Ref. Prosen ; Prosen07 ) that there are large section of $\tau$ values where $|\lambda_{1}|$ appears to be essentially $1$, which would suggest that correlations do not decay, i.e., as if one would have an almost “integrable” system (rigorous integrability is of course highly unlikely). Note that we do not have in mind the interval $1.5\lesssim\tau\lesssim 2.0$ where one indeed has three quite close integrable points at $\tau=\pi/2$ (integrability type (iv) from the introduction), and $\tau=\pi/(2h_{\rm x})$ and $\tau=\pi/(2h_{\rm z})$ (type (iii) integrability), but regions outside of this interval. However, upon inspecting other sectors we see that things are different. The values of $|\lambda_{1}|$ there are smaller than $1$, suggesting decay of correlations. We shall explore in more details all those situations for the three indicated selected values of $\tau$ in the following sections. For now let us just say that they correspond to different dynamical regimes, e.g., chaotic with exponential mixing vs. just mixing, and that having access to those other momentum sectors is instrumental in determining the correct dynamical behavior.
Finally, let us comment on the convergence of $\lambda_{1}$ with $r$. We can see in Fig. 2 that the speed of convergence varies very much with $\tau$ as well as with $k$. For instance, in the $k=0^{+}$ sector $|\lambda_{1}|$ has converged already at $r=10$ for almost all shown values of $\tau$ (the red for $r=10$ and the orange curve for $r=8$ are indistinguishable except in a very narrow region around the minimum at $\tau\approx 0.8$). For $k=0^{-}$ the convergence is good for $\tau>0.65$, but much worse for smaller $\tau$. The situation is least converged for $k=\pi$ (note also that we show only $r=8$), however, there is again a strong asymmetry; the eigenvalue is converged for larger $\tau$, while not at smaller $\tau$. Slow convergence at very small $\tau$ can likely be traced back to the fact that all eigenvalues are very close to $1$, e.g., both the inner and outer radius (18) of a ring that contains most of the eigenvalues converge to $1$ as $\tau\to 0$. For $\tau$ that is an even multiple of $\pi/4$ (the dips in Fig. 2) the inner radius $r_{\rm in}$ instead becomes 00, however the number of nonzero eigenvalues of $M(k)$ is smaller and the matrix is strongly non-normal, likely contributing to slow convergence with $r$.
The rich dependence on three relevant parameters calls for a detailed investigation of the KI phase diagram. In the present work we will focus only on three representative cases, showing that there are different dynamical regimes in a quantum many body system. In particular, as we shall see, the system is not always chaotic (see also previous results in Ref. Prosen ).
@@ III Quantum chaotic regime
Considering that all $|\lambda_{1}(k)|<1$ we can say that we are in a quantum chaotic regime. Therefore one expects that all correlation functions should asymptotically decay to zero exponentially as $\sim|\lambda_{1}|^{t}$, and, considering that the eigenvalue depends on $k$, different operators will decay with a different rate. We show in Fig. 4 that this is indeed the case. We compare numerically calculated $C(t)$ and the predicted asymptotic decay based on the leading RP resonance, always finding agreement with values in Fig. 3.
@@
decay the slowest. On the other hand, the “current” $J$ (12) decays faster because it is from the odd $k=0^{-}$ sector that has smaller $|\lambda_{1}|$. Still faster is the decay of the staggered magnetization $sZ$ (13), which is from $k=\pi$ ($sZ$ decays in an oscillatory manner therefore in this case we always plot $|C(t)|$). By superimposing extensive operators with different $k$ we can study operators with any spatial dependence. The extreme case is a strictly local operator, for instance polarization $\sigma^{\rm x}_{L/2}$ at the middle site, which is a uniform superposition of all $k$ modes. Because the $k=0^{+}$ contribution will decay the slowest we expect that such a local autocorrelation function will asymptotically also decay as $|\lambda_{1}(k=0^{+})|^{t}$. The correlation function, defined in this case without an $1/L$ prefactor,
@@ III.1 Slow convergence with system size
However, looking at the magnetization $Z$ (20) in Fig. 7(a) an exponential decay is seen with a rate that is not given by $|\lambda_{1}(0^{+})|$. This is rather puzzling considering that the RP resonance $\lambda_{1}(0^{+})$ converges rather quickly (Fig. 5), and that this seeming incorrect decay persists over almost 3 decades all the way till finite-size fluctuations become noticeable at very small values of $C(t)$. To understand how that can be we have looked at the evolution of $Z$ by the truncated propagator $M(0^{+})$ whose largest eigenvalue is clearly $\lambda_{1}\approx 0.825$ (and is different from the observed decay that goes as $\sim 0.76^{t}$). Namely, within the truncation to $r$-site operators the corresponding correlation function can be simply written as
@@
where the initial vector $\mathbf{y}$ corresponds to the operator $a\in{\cal S}_{r}$, in our case to $a=\sigma^{\rm z}$. In a basis where $a=\sigma^{\rm z}$ is chosen as one basis vector one therefore needs only the appropriate diagonal matrix element of the truncated matrix $M^{t}$.
@@
In Fig. 7(b) we can see that the above truncation approximation to $C(t)$ (Eq.24, curves with small symbols in Fig. 7(b)) does asymptotically decay as $\lambda_{1}^{t}$, but the asymptotic decay starts rather late, say around $t>40$ when $C(t)$ is already quite small (because the corresponding expansion coefficient is small). We can also see that, while the $r=6$ approximation does not yet capture the true decay for $t<30$, higher truncations like $r=10$ or $r=12$ do. We can conclude that in order to see the true asymptotic decay $|\lambda_{1}|^{t}$ one needs to look at small values $C(t)<10^{-6}$, and therefore in order to even get a glimpse of it in a single random state evolution one would need large systems with $L>40$.
Such decay, where one initially has faster relaxation $\sim 0.76^{t}$ that then eventually goes into asymptotic $\sim 0.82^{t}$ given by the largest eigenvalue, is reminiscent of a recently discovered two-step relaxation in random circuits PRX in which the initial decay is determined by the pseudospectrum rather than the spectrum U4 . It can happen when dealing with non-normal matrices, e.g., non-Hermitian, that are badly conditioned. In such cases the spectrum is a rather singular object that is very sensitive to perturbations while the pseudospectrum trefethen is robust. It seems though that the effect here is different in one crucial aspect: the window in which one has the “incorrect” decay $\sim 0.76^{t}$ is here of fixed width whereas in the mentioned random circuits it diverges with system size. This divergence comes because the condition number of those matrices in random circuits diverges with system size, allowing for an increasing window size as one increases the system size. To get an idea on the condition number of our $M(k=0^{+})$ we have looked at its singular values. Namely, for a nonsingular matrix $M$ its condition number is simply $\kappa=s_{\rm max}(M)/s_{\rm min}(M)$, where $s_{\rm max}$ and $s_{\rm min}$ are maximal and minimal singular values. Using exact expressions for singular values (14) we see that the the maximal one is $1$, while the minimal one is $\cos^{2}{(2\tau)}$, both independent of $r$ or any other parameter. The matrix $M$ becomes singular only at special $\tau=\pi/4\approx 0.78$. Because our $\tau=0.75$ is very close to that singularity we have a rather large $\kappa\approx 200$, which though is finite and in particular does not increase with $r$. Therefore, there is no true two-step relaxation; the initial decay is just a transient that happens to hold for a rather large range of $C(t)$. Nevertheless, slow convergence of the autocorrelation function towards its asymptotic decay can be traced back to strong non-normality of $M(k)$. An interesting open question is, whether perhaps at $\tau=\pi/4$ one could have some effects of strong non-normality due to some singular values becoming zero.